OpenStudy (anonymous):

F(x,y) =sinx + siny + sin(x+ y)  find the critical points of the local max and min im kinda stuck on this one, please help?

OpenStudy (experimentx):

to find the critical point you need to equate first derivative to zero.

OpenStudy (anonymous):

both partials need to be 0

OpenStudy (experimentx):

do it simultaneously ... you'll have a system of two equations in x and y. solve for x and y

OpenStudy (anonymous):

well yeah i knw how to do that already fx=cosx+cos(x+y) fy=cosy+cos(x+y)

OpenStudy (anonymous):

idk how to solve for x and y tho

OpenStudy (experimentx):

so what are the critical points?

OpenStudy (anonymous):

thats wht im looking for

OpenStudy (anonymous):

This is a function of x and y, it's a surface in R3

OpenStudy (anonymous):

LOL i dont need u to explain that to me i knw that already i just need to knw how to solve for x and y to find the critical points

OpenStudy (anonymous):

if you knw how to do that could u please tell me? cuz it seems difficult to isolate X or Y by its self with this function

OpenStudy (turingtest):

perhaps we can use cramer's rule here?

OpenStudy (anonymous):

i forgot how to do that

OpenStudy (experimentx):

hmm ... what do we get here?  use cos D + cos C = 0